OpenStudy (a1234):

Use an inverse trigonometric function to write theta as a function of x. I'll draw...

OpenStudy (a1234):



OpenStudy (a1234):

Is it theta = arctan((x+1)/10)?

Nnesha (nnesha):

yes right \[\rm sin \rm \theta = \frac{ opposite }{ hypotenuse }~~~~ \cos \theta = \frac{ adjacent }{ hypotenuse } ~~\tan \theta = \frac{ opposite }{ adjacent }\] tan =opposite/adjacent